OpenStudy (anonymous):

Could someone give me an example of how to solve these types of problems? Using the following information calculate the test statistic z. ( 2 points each) a.) Ho: µ = 10, σ = 3, n = 40, = 10.6 b.) Ho: µ = 120, σ = 23, n = 25, = 126.2 c.) Ho: µ = 18.2, σ = 3.7, n = 140, = 18.93 d.) Ho: µ = 81, σ = 13.3, n = 50, = 79.6

OpenStudy (anonymous):

3+40=43 so add 10.6 10.6 rounds to 11 to 43+11=

OpenStudy (anonymous):

i had this question and that was the answer

OpenStudy (anonymous):

So the answer would be 54?

OpenStudy (anonymous):

For B, I would then do 23+25= 48 then add 126.2?

ganeshie8 (ganeshie8):

test statistic, z = (observation - "null hypothesis mean")/(standard error)

ganeshie8 (ganeshie8):

start by finding the standard error

ganeshie8 (ganeshie8):

a.) Ho: µ = 10, σ = 3, n = 40, = 10.6 observation = 10.6 null hypothesis mean = 10

ganeshie8 (ganeshie8):

find out "standard error" and plug it in the earlier formula

OpenStudy (anonymous):

How do I solve standard error?

ganeshie8 (ganeshie8):

standard error = \(\large \frac{\sigma}{\sqrt{n}}\)

OpenStudy (anonymous):

So for the first one, 3/square root of 40?

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

1) Standard error = 3/square root of 40 = 0.47 10.6-10 / 0.47?

ganeshie8 (ganeshie8):

z = (observation - "null hypothesis mean")/(standard error) = \(\large \frac{10.6 - 10}{\frac{3}{\sqrt{40}}}\)

ganeshie8 (ganeshie8):

yes, simplify.. u may work out all others same way

OpenStudy (anonymous):

.6/0.47 = 1.27 as the final answer?

ganeshie8 (ganeshie8):

correct !

OpenStudy (anonymous):

Thank you!

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

?